# Self-gravitating solutions in Yang-Mills–Chern-Simons theory coupled to 3D massive gravity

### v1

@@ IV Warped AdS black hole
Having studied the case of the 3-sphere and the squashed 3-sphere, we are ready to present the solution for the WAdS3 black hole. As explained in the introduction, the argument goes as follows: AdS3 space can be written as Hopf fibration, just like in the case of the 3-sphere, by simply replacing trigonometric functions in (7) by their hyperbolic counterparts. Then, the WAdS3 spaces are nothing but the $\alpha$-deformations of AdS3, exactly as done in (14) for the squashed 3-sphere. Having solved the case of positive curvature, the hyperbolic analog can easily be obtained by continuing the gauge theory to the $SU(1,1)$ group. Then, taking into account the observation of Anninos et al. (2009), which states that the WAdS3 black holes of Moussa et al. (2008) are discrete quotients of WAdS3, we are ready to derive the solutions we are looking for. Following the same steps as before, but adapting them to the hyperbolic case, we are led to find the adequate relations between the coupling constants of the theory and the squashed parameter $\alpha$. We also take $k=0$. Then, we can consider the stationary ansatz
together with the following ansatz for the gauge field
$A=\lambda^{ij}\omega_{i}\eta_{j}\,,$ (22)
where $\lambda^{ij}=\operatorname{diag}(\lambda_{(1)},\lambda_{(2)},\lambda_{(3)})$ is in the adjoint representation of $SU(1,1)$; $\lambda_{(i)}$ are constants, and $\omega_{i}$ are the left-invariant forms of $SU(1,1)$. As it happens for NMG in vacuum Clément (2009), the field equations (3) are solved by the following functions
$f(r)$ $=\frac{\ell^{2}(r-r_{+})(r-r_{-})}{r\left[2\alpha\sqrt{r_{+}r_{-}}+(\alpha^{2}-1)r+r_{+}+r_{-}\right]}\,,$ (23)
$h(r)$ $=\frac{(r-r_{+})(r-r_{-})}{\rho_{0}^{2}}\,,$ (24)
$\rho^{2}(r)$ $=\frac{r\ell^{4}\left[2\alpha\sqrt{r_{+}r_{-}}+(\alpha^{2}-1)r+r_{+}+r_{-}\right]}{4\rho_{0}^{2}}\,,$ (25)
$N^{\theta}(r)$ $=\frac{2\rho_{0}\left[\sqrt{r_{+}r_{-}}+\alpha\,r\right]}{r\ell\left[2\alpha\sqrt{r_{+}r_{-}}+(\alpha^{2}-1)r+r_{+}+r_{-}\right]}\,.$ (26)
where $\Lambda=-\ell^{-2}$. $r_{\pm}$ are two integration constants that, provided $0\leq r_{-}\leq r_{+}$, represent the locations of the outer and the inner horizons of the black hole, cf. Anninos et al. (2009). We also find that the following relations must be obeyed
$\rho_{0}^{2}$ $=\frac{\left[\lambda_{(2)}-\lambda_{(1)}\lambda_{(2)}\right]^{2}-\alpha^{2}\big{[}\lambda_{(1)}-\lambda_{(2)}^{2}\big{]}^{2}}{2\kappa e^{2}\alpha^{2}(\alpha^{2}-1)}+\frac{21\alpha^{2}-4}{8\mu}\,,$ (27)
$\frac{1}{\ell^{2}}$ $=-\frac{1}{\rho_{0}^{4}}\Bigg{[}\frac{21\alpha^{4}-72\alpha^{2}+16}{64\mu}$
$+\frac{\alpha\left(9\alpha^{4}-17\alpha^{2}-24\right)\sqrt{9\alpha^{2}+16}-27\alpha^{4}\left(\alpha^{2}-1\right)+128\alpha^{2}+32}{128e^{2}\kappa\left(\alpha^{2}-1\right)}\Bigg{]}\,,$ (28)
$\lambda_{(1)}$ $=1-2\lambda_{(2)}^{2}\,,$ (29)
$\lambda_{(2)}^{2}$ $=\lambda_{(3)}^{2}=\frac{\alpha}{8}\left(\sqrt{9\alpha^{2}+16}-3\alpha\right)\,.$ (30)
Notice that all these parameters get fixed in terms of $\alpha$ which, in turn, is a function of the AdS3 radius through (28).
The field configuration (22) can be seen to be finite at the horizon and to exhibit an asymptotic behavior similar to the one obtained in Bañados et al. (2006) for the $U(1)$ gauge field in asymptotically WAdS3 spaces.
The conserved charges associated with the WAdS3 black holes have been computed in Donnay and Giribet (2015); see also references therein and thereof. A similar computation works here.
We can also parametrize the left-invariant forms of $SU(1,1)$ as follows
$\omega_{1}$ $=\text{d}u+\sinh\sigma\text{d}\tau\,,$ (31a)
@@
$\omega_{3}$ $=\sinh u\text{d}\sigma-\cosh u\cosh\sigma\text{d}\tau\,,$ (31c)
with $(u,\sigma,\tau)$ being scalar functions of the coordinates $(t,r,\theta)$; they are explicitly given by
@@
which are defined for the range $21\leq\alpha<$ 2 Anninos et al. (2009). Coordinates $(u,\sigma,\tau)$ are useful to describe locally WAdS3 spaces as fibrations and explore their properties.
Notice that the limit $\alpha\to 1$ of the solution above is smooth and yields $\lambda_{(i)}=1/2$ for $i=1,2,3$. In this limit, we have
$\rho_{0}^{2}=\frac{85\kappa e^{2}+\mu}{40\kappa e^{2}\mu}\;\;\;\;\;\mbox{and}\;\;\;\;\;\frac{1}{\ell^{2}}=\frac{35\kappa e^{2}\mu(25\kappa e^{2}+\mu)}{(85\kappa e^{2}+\mu)^{2}}\,.$ (33)
This is the limit in which the squashing effect disappears and the solution is locally equivalent to AdS3. Another limit that is worth looking at is $e\to\infty$, where the solution above is seen to agree with the WAdS3 black hole of NMG in vacuum, originally found in Ref. Clément (2009).
For arbitrary values of the parameters, the solution (23)-(30) yields a non-vanishing stress-energy tensor, $T_{\mu\nu}$, which can be written as a linear combination
@@
which can be seen to be finite in the limit $\alpha\to 1$. Equation (34) implies that the YM field shifts the cosmological constant $\Lambda$ and the mass parameter $\mu$. From the expressions for $c_{1}$ and $c_{2}$ one can verify that the WAdS3 solution may exist even if the higher-curvature terms of NMG are not present, i.e. in the limit $\mu\to\infty$. That is to say, the YM terms suffice to support the black hole background in Einstein gravity. In some cases, the meron-type ansatz renders the YM configuration locally equivalent to an Abelian gauge field Canfora et al. (2022); in the case of the WAdS3 black holes this seems to be different: while the WAdS3 black holes are known to exist in the Einstein gravity coupled to the $U(1)$ gauge field theory Bañados et al. (2006), in that case, the CS term is needed to support the background. It is also worth mentioning that the solution we derived here persists when the gravitational CS term of Topologically Massive Gravity is also included. This merely induces a shift in the relation among the coupling constants of the theory and the parameters of the solution.
V Other solutions and Final remarks
@@
Here, we have constructed exact black hole solutions of 3-dimensional massive gravity sourced by a non-Abelian gauge field. The solutions asymptote WAdS3 at large distance and coincide with the NMG solution in vacuum found in Clément (2009) in the appropriate limit. This result might be relevant to the study of WAdS3/(W)CFT2 correspondence Anninos et al. (2009); Detournay et al. (2012); Donnay and Giribet (2015, 2017) and it would be interesting to study the implications in that context. Among the solutions we found, there are self-gravitating black holes in Einstein–Yang-Mills theory.
@@
The theory we considered throughout the paper was NMG coupled to YM-CS gauge theory. This theory is of fourth order and is non-linear both in the gravitational and the matter sector, which makes the problem of solving the field equations analytically quite intricate. This demanded from us to find an adequate ansatz that resulted in a simplification of the equations. We considered a meron-type ansatz for the gauge field, which renders the problem tractable analytically. This ansatz turns out to be particularly useful to find exact solutions to this theory. In order to illustrate that and to show how the results of this paper can be extended, let us present some additional cases below.
@@
One relatively simple example is the Nil geometry, which is one of the eight Thurston geometries that appear in connection to the geometrization conjecture. It can be associated to a Bianchi II space endowed with a transitive nilpotent group of diffeomorphisms. Locally, it remains invariant under the action of the Heisenberg group; see Refs. Gegenberg and Kunstatter (2004); Chernicoff et al. (2018); Flores-Alfonso et al. (2021b). We write its Euclidean metric as
@@
where we denote $\rho$ as the radius.
There are two, non-flat, left-invariant metrics on the Nil group with Lorentzian signature Rahmani (1992). One of them corresponds to fiber over Euclidean space,
@@
and the other is a fiber over Lorentzian space,
@@
They are sometimes referred to as warped flat spaces Anninos (2010). As pointed out in Ref. Mukhopadhyay et al. (2014), the first of these metrics admits the interpretation of a cross-section of the cylindrical Som-Raychauduri spacetimes Som and Raychaudhuri (1968). On the other hand, metric (38) can be interpreted as a Kundt spacetime, a natural generalization of the $pp$-waves; see Chow et al. (2010); Boucetta et al. (2022); Flores-Alfonso et al. (2023).
Using an ansatz for the gauge field similar to the one used in the previous sections, we can show that NMG coupled to YN-CS theory also admits Nil manifolds as solutions. To show this, let us focus on the solution (37) and take the $SU(2)$ gauge field
@@
$\lambda$ $=\frac{1}{2}\,,$ $\rho^{2}$ $=\frac{1}{4k^{2}}\,,$ $\Lambda$ $=\frac{k^{2}\left(\kappa e^{2}+2k^{2}\right)}{2\kappa e^{2}}\,,$ $\mu$ $=\frac{21\kappa e^{2}k^{2}}{2\left(\kappa e^{2}-2k^{2}\right)}\,,$ (40)
@@
Another interesting type of solutions of NMG that support non-trivial gauge field configuration is the AdS-waves, which are analogs of $pp$-waves but in AdS3, cf. Ayón-Beato and Hassaïne (2006); Ayón-Beato et al. (2009b). The metric os these solutions takes the form
@@
where $u\in\mathbb{R}$ and $v\in\mathbb{R}$ are null coordinates, $\rho\in\mathbb{R}_{+}$ is a constant, and $F(u,y)$ is a function that represents the profile of the wave and is obtained by solving the field equations. Locally AdS3 corresponds to the particular case $F=\text{const}$. In these coordinates, we consider a convenient parameterization of the $SU(1,1)$ gauge field, namely
$A=$ $2\lambda\Bigg{[}\left(\frac{2(v-1)}{y^{2}}{\rm d}u-\frac{1}{y}{\rm d}y\right)\eta_{1}+\left(\frac{v(v-2)}{y^{2}}{\rm d}u+\frac{{\rm d}v}{2}-\frac{v-1}{y}{\rm d}y\right)\eta_{2}$
$\quad-\left(\frac{v^{2}-2v+2}{y^{2}}{\rm d}u+\frac{{\rm d}v}{2}-\frac{v-1}{y}{\rm d}y\right)\eta_{3}\Bigg{]}\,.$ (42)
@@
where $c(u)$ is an arbitrary function of the null coordinate $u$. Inserting these solutions into the remaining field equations, we find that they are solved by
$\rho^{2}$ $=\frac{\kappa e^{2}-2\mu}{2\left[\kappa e^{2}\left(k^{2}-\mu\right)-k^{2}\mu\right]}\,,$ (44a)
$\Lambda$ $=-\frac{4\mu^{3}\left(5\kappa e^{2}+4k^{2}\right)+4\mu^{2}\left(k^{4}-5\kappa e^{2}k^{2}-\kappa^{2}e^{4}\right)-7\mu\kappa e^{2}k^{4}+4\kappa^{2}e^{4}k^{4}}{4\mu\left(\kappa e^{2}-2\mu\right)^{2}}\,.$ (44b)
@@
Acknowledgements.
The authors thank Fabrizio Canfora for discussions. The work of C.C. is supported by Agencia Nacional de Investigación y Desarrollo (ANID) through FONDECYT Regular grants No 1240043, 1240048, 1230112, and 1210500. D.F-A. is supported by ANID under FONDECYT grant No. 3220083. J.O. acknowledges the financial support of FONDECYT Regular grant No 1221504.

### v2

@@ IV Warped AdS black hole
Having studied the case of the 3-sphere and the squashed 3-sphere, we are ready to present the solution for the WAdS3 black hole. As explained in the introduction, the argument goes as follows: AdS3 space can be written as a Hopf fibration over AdS2, just like in the case of the 3-sphere, by simply replacing trigonometric functions in Eq. (7) by their hyperbolic counterparts. Then, the WAdS3 spaces are nothing but the $\alpha$-deformations of AdS3, exactly as done in (14) for the squashed 3-sphere. Having solved the case of positive curvature, the hyperbolic analog can easily be obtained by continuing the gauge theory to the $SU(1,1)$ group; the latter is isomorphic to $SL(2,\mathbb{R})$. For further details on how this process is carried out see Sec. 4 of Duff et al. (1999).
We begin by writing warped AdS3 space in analogy with (14), i.e.,
$\text{d}s^{2}=\rho_{0}^{2}\left[-\cosh^{2}\sigma\text{d}\tau^{2}+\text{d}\sigma^{2}+\alpha^{2}\left(\text{d}u+\sinh\sigma\text{d}\tau\right)^{2}\right]\,.$ (20)
At the same time, we require the left-invariant forms of the gauge group to construct a gauge potential that serves our purposes. In the hyperbolic coordinates used above, they can be parametrized as
$\omega_{1}$ $=-\sinh u\text{d}\sigma+\cosh u\cosh\sigma\text{d}\tau\,,$ (21a)
@@
$\omega_{3}$ $=\text{d}u+\sinh\sigma\text{d}\tau\,.$ (21c)
These $1$-forms satisfy the Maurer-Cartan equations $\text{d}\omega_{i}+\tfrac{1}{2}f_{ijk}\omega^{j}\wedge\omega^{k}=0$, with group indices being raised and lowered with the Cartan-Killing metric of $SU(1,1)$, say $\eta_{kl}=\operatorname{diag}(-1,1,1)$, with $f_{ijk}=\epsilon_{ijl}\eta_{kl}$ the structure constants thereof. Moreover, in terms of them, the spacetime metric takes the form $\text{d}s^{2}=\rho_{0}^{2}\;\left(-\omega_{1}^{2}+\omega_{2}^{2}+\alpha^{2}\omega_{3}^{2}\right)$, bearing a close resemblance with that of Eq. (7). For the non-Abelian gauge fields, we assume the following ansatz
$A=\sum_{i=1}^{3}\lambda_{(i)}\omega_{i}\eta_{i}\,,$ (22)
where the $\lambda_{(i)}$ are constants, and $\eta_{i}$ are the generators of the gauge group satisfying the ${su}(1,1)$ Lie algebra $[\eta_{i},\eta_{j}]=f_{ijk}\eta_{k}$.
Following the same steps as before, but adapting them to the hyperbolic case, we are led to find the adequate relations between the coupling constants of the theory and the squashing parameter $\alpha$. We also take $k=0$, as in the previous section, to avoid higher-degree polynomials. Then, the field equations of NMG coupled to Yang-Mills fields are solved by
$\rho_{0}^{2}$ $=\frac{(\lambda_{(2)}-\lambda_{(2)}\lambda_{(3)})^{2}-\alpha^{2}(\lambda_{(2)}^{2}-\lambda_{(3)})^{2}}{2\kappa e^{2}\alpha^{2}(\alpha^{2}-1)}-\frac{4-21\alpha^{2}}{8\mu^{2}}\,,$ (23a)
$\Lambda$ $=-\frac{\alpha^{2}}{4\rho_{0}^{2}}+\frac{63\alpha^{4}-80\alpha^{2}+16}{64\mu^{2}\rho_{0}^{4}}-\frac{(\lambda_{(2)}^{2}-\lambda_{(3)})^{2}}{4\kappa e^{2}\rho_{0}^{4}}\,,$ (23b)
while the $\lambda_{(i)}$ parameters satisfy the same conditions as the squashed sphere, Eq. (18). All the parameters get fixed in terms of $\alpha$ which, in turn, is a function of the cosmological constant. Notice that this solution can be obtained from Eq. (19) by performing $\mu^{2}\to-\mu^{2}$ and $\Lambda\to-\Lambda$.
Now, we take into account the observation of Ref. Anninos et al. (2009), which states that the WAdS3 black holes of Moussa et al. (2008) are discrete quotients of WAdS3. Thus, we are ready to derive the solutions we are looking for. First, we perform the change of coordinates
@@
which are defined for the range $21\leq\alpha<$ 2 (see Anninos et al. (2009)). Then, as it happens for NMG in vacuum Clément (2009), one finds that the field equations (3) are solved by the stationary metric
with the metric functions
$f(r)$ $=\frac{\ell^{2}(r-r_{+})(r-r_{-})}{r\left[2\alpha\sqrt{r_{+}r_{-}}+(\alpha^{2}-1)r+r_{+}+r_{-}\right]}\,,$ (26a)
$h(r)$ $=\frac{(r-r_{+})(r-r_{-})}{\rho_{0}^{2}}\,,$ (26b)
$\rho^{2}(r)$ $=\frac{r\ell^{4}\left[2\alpha\sqrt{r_{+}r_{-}}+(\alpha^{2}-1)r+r_{+}+r_{-}\right]}{4\rho_{0}^{2}}\,,$ (26c)
$N^{\theta}(r)$ $=\frac{2\rho_{0}\left[\sqrt{r_{+}r_{-}}+\alpha\,r\right]}{r\ell\left[2\alpha\sqrt{r_{+}r_{-}}+(\alpha^{2}-1)r+r_{+}+r_{-}\right]}\,,$ (26d)
where $\Lambda=-\ell^{-2}$. The two integration constants $r_{\pm}$ represent the locations of the outer and the inner horizons of the black hole provided $0\leq r_{-}\leq r_{+}$, cf. Anninos et al. (2009).
The field configuration (22) can be seen to be finite at the horizon and to exhibit an asymptotic behavior similar to the one obtained in Bañados et al. (2006) for the $U(1)$ gauge field in asymptotically WAdS3 spaces. Additionally, the conserved charges associated with the WAdS3 black holes have been computed in Donnay and Giribet (2015); see also references therein and thereof. A similar computation works here, as the asymptotic behavior of the Yang-Mills fields translates into a shift of $\rho_{0}$ and the WAdS3 curvature radius of the black hole. Then, in our notation, the mass and angular momentum of the WAdS3 black hole are Donnay and Giribet (2015)
$Q[\partial_{t}]$ $:=M=\frac{6\alpha}{\ell G(4-\alpha^{2})(4-21\alpha^{2})}\left[2\sqrt{r_{+}r_{-}}-\alpha(r_{+}+r_{-})\right]\,,$ (27)
$Q[\partial_{\theta}]$ $:=J=\frac{6\sqrt{3}\,\alpha}{2\ell^{2}\mu^{2}G(1-\alpha^{2})(4-\alpha^{2})^{3/2}}\left[\alpha\sqrt{r_{+}r_{-}}(r_{+}+r_{-})-r_{+}r_{-}(\alpha^{2}+1)\right]\,,$ (28)
respectively, where $\alpha$ and $\ell$ are related by Eq. (23b). In the limit $\alpha\to 1$, the squashing effect disappears and the solution is locally equivalent to AdS3. For this limiting case the solution above smoothly yields $\lambda_{(i)}=1/2$ for $i=1,2,3$, and thus we have
$\rho_{0}^{2}=\frac{85\kappa e^{2}+\mu^{2}}{40\kappa e^{2}\mu^{2}}\;\;\;\;\;\mbox{and}\;\;\;\;\;\frac{1}{\ell^{2}}=\frac{35\kappa e^{2}\mu^{2}(25\kappa e^{2}+\mu^{2})}{(85\kappa e^{2}+\mu^{2})^{2}}\,.$ (29)
Another limit that is worth looking at is $e\to\infty$, where the solution above is seen to agree with the WAdS3 black hole of NMG in vacuum, originally found in Ref. Clément (2009).
Notice that the black hole is $\eta$-Einstein with $\eta=\omega_{3}$. Thus, the NMG tensor satisfies a similar decomposition as Eq. (15). This geometric property ultimately informs the energy-momentum tensor via the equations of motion. Hence, for arbitrary values of the parameters, the solution (23) yields a nonvanishing stress-energy tensor, $T_{\mu\nu}$, which can be written as a linear combination
@@
being finite as $\alpha\to 1$. Equation (30) implies that the YM field shifts the cosmological constant $\Lambda$ and the inverse of the mass parameter $\mu$. From the expressions for $c_{1}$ and $c_{2}$ one can verify that the WAdS3 solution may exist even if the higher-curvature terms of NMG are not present, i.e. in the limit $\mu\to\infty$. That is to say, the YM terms suffice to support the black hole background in Einstein gravity, even when $\Lambda=0$. In some cases, the meron-type ansatz renders the YM configuration locally equivalent to an Abelian gauge field Canfora et al. (2022); in the case of the WAdS3 black holes this seems to be different: while the WAdS3 black holes are known to exist in the Einstein gravity coupled to the $U(1)$ gauge field theory Bañados et al. (2006), in that case, the CS term is needed to support the background. It is also worth mentioning that the solution we derived here persists when the gravitational CS term of Topologically Massive Gravity is also included. This merely induces a shift in the relation between the coupling constants of the theory and the parameters of the solution.
V Final Remarks
@@
In this work, we constructed exact warped-AdS black hole solutions in NMG minimally coupled to a non-Abelian gauge field. The solution asymptotes WAdS3 at large radial distances and coincides with the NMG solution in vacuum found in Clément (2009) in the appropriate limit. This result might be relevant to the study of WAdS3/WCFT2 correspondence Anninos et al. (2009); Detournay et al. (2012); Donnay and Giribet (2015, 2017), as rotating black hole solutions sourced by self-gravitating nonlinear matter fields are scarce.
@@
The theory under consideration is of fourth order and is non-linear both in the gravitational and the matter sector, which makes the problem of solving the field equations analytically quite intricate. Our strategy exploits the construction of $3$-dimensional spacetimes as Hopf vibrations of $2$-dimensional ones, resulting in a simplification of the field equations. We considered a meron-type ansatz for the gauge field, which renders the problem tractable analytically. This ansatz turns out to be particularly useful to find exact solutions to this theory. Indeed, the solution can be perfectly embedded in Einstein-Yang-Mills theory as it is continuously connected to the latter when the higher-curvature terms are absent.
Interesting questions remain open. For instance, in the AdS/CFT correspondence, holographic DC conductivities can be obtained from planar AdS black holes with Maxwell sources Donos and Gauntlett (2014); Andrade and Withers (2014); Caldarelli et al. (2017). Holographic responses of Yang-Mills sources are worth exploring in chiral theories using the WAdS3/WCFT2 correspondence, as they could provide analytic expressions for the vacuum expectation value of strongly coupled non-Abelian currents. Additionally, null WAdS3 spacetimes, also discussed in Ref. Anninos et al. (2009), are known to be diffeomorphic to time-independent AdS waves observing the Schrödinger symmetry. Thus, looking for solutions of this type with self-gravitating Yang-Mills fields becomes relevant for describing condensed matter system holographically Son (2008); Balasubramanian and McGreevy (2008). Finally, it is known that squashed 3-spheres contribute nontrivially to the parity anomaly in three dimensions Redlich (1984a, b); Alvarez-Gaume et al. (1985). As this space bears a close resemblance to WAdS3, a similar effect could be induced by the solution found here. We postpone a deeper study of these questions for future works.
Acknowledgements.
The authors thank Fabrizio Canfora and Rodrigo Olea for discussions. This work is supported by Agencia Nacional de Investigación y Desarrollo (ANID) through FONDECYT grants No 1240043, 1240048, 1230112, 1210500, 1221504, and 3220083.
Appendix A Other Solutions
In order to further illustrate the usefulness of meron-type ansätze in this context, we provide explicit examples which extend the results presented above. We begin by returning to the squashed 3-sphere. Let us mention that the field ansatz (8) is related to a geometric mapping of a three-sphere in spacetime into the three-sphere underlying the $SU(2)$ gauge group. Since the squashed three-sphere fibers over a two-sphere, a natural alternative is to consider a map from the sphere into the equator of the gauge group. This approach yields the following gauge potential
$A$ $=2\lambda\,\Big{[}\left(\sin\varphi\,{\rm d}\vartheta+\cos\varphi\cos\vartheta\sin\vartheta\,{\rm d}\varphi\right)t_{1}$
$\qquad-\left(\cos\varphi\,{\rm d}\vartheta-\sin\varphi\cos\vartheta\sin\vartheta\,{\rm d}\varphi\right)t_{2}-\sin^{2}\vartheta\,{\rm d}\varphi\,t_{3}\Big{]}\,.$ (32)
This choice, however, leads to a gauge field with a trivial topological charge, as the gauge connection depends only on two coordinates of the squashed 3-sphere. Inserting Eq. (14) and the ansatz (32) into the field equations (3), we find that the latter reduces to a system of algebraic equations that admit the following solution
$\lambda=\frac{1}{2}\;,\;\;\;\;\;\;\;\;\rho^{2}=\frac{\alpha^{2}}{4k^{2}}\,,$ (33)
with the coupling constants being related as follows
$\mu^{2}$ $=-\frac{k^{2}\kappa\,e^{2}\left(21\alpha^{4}-25\alpha^{2}+4\right)}{2\alpha^{4}e^{2}\kappa-2\alpha^{2}e^{2}\kappa+4k^{2}}\,.$ (35)
A real, Lorentzian version of the solution can also be obtained by performing the analytic continuation $\alpha\to i\alpha$. The relevance of solutions of this type is that they circumvent the need for anisotropy in the gauge field.
@@ Appendix A Other Solutions
A similar example we consider is the Nil geometry, which is one of the eight Thurston geometries that appear in connection to the geometrization conjecture. It can be associated to a Bianchi II space endowed with a transitive nilpotent group of diffeomorphisms. Locally, it remains invariant under the action of the Heisenberg group; see Refs. Gegenberg and Kunstatter (2004); Chernicoff et al. (2018); Flores-Alfonso et al. (2021b). We write its Euclidean metric as
@@
where we denote $\rho$ as the radius. There are two, non-flat, left-invariant metrics on the Nil group with Lorentzian signature Rahmani (1992). One of them has a base space which is the Euclidean plane,
@@
and the other is a fiber bundle over the Minkowski plane,
@@
They are sometimes referred to as warped flat spaces Anninos (2010). As pointed out in Ref. Mukhopadhyay et al. (2014), the first of these metrics admits the interpretation of a cross-section of the cylindrical Som-Raychauduri spacetimes Som and Raychaudhuri (1968). On the other hand, metric (38) can be interpreted as a Kundt spacetime, a natural generalization of the $pp$-waves; see Chow et al. (2010); Boucetta et al. (2022); Flores-Alfonso et al. (2024). Using a geometrically similar ansatz for the gauge field as (32), we can show that NMG coupled to YM-CS theory also admits Nil manifolds as solutions. To show this, let us focus on the solution (37) and take the $SU(2)$ gauge field
@@
$\lambda$ $=\frac{1}{2}\,,$ $\rho^{2}$ $=\frac{1}{4k^{2}}\,,$ $\Lambda$ $=\frac{k^{2}\left(\kappa e^{2}+2k^{2}\right)}{2\kappa e^{2}}\,,$ $\mu^{2}$ $=\frac{21\kappa e^{2}k^{2}}{2\left(\kappa e^{2}-2k^{2}\right)}\,,$ (40)
@@
Other interesting solutions of NMG that support a non-trivial gauge field configuration are AdS-waves, which are analogs of $pp$-waves but in AdS3, cf. Ayón-Beato and Hassaïne (2006); Ayón-Beato et al. (2009b). The metric of these solutions takes the form
@@
where $u\in\mathbb{R}$ and $v\in\mathbb{R}$ are null coordinates, $\rho\in\mathbb{R}_{+}$ is a constant, and $F(u,y)$ is a function that represents the profile of the wave and is obtained by solving the field equations. Locally AdS3 corresponds to the particular case $F=\text{const}$. In these coordinates, the $SU(1,1)$ gauge field $A=\lambda\omega_{i}\eta_{i}$ is parameterized as
$A=$ $2\lambda\Bigg{[}\left(\frac{v^{2}-2v+2}{y^{2}}{\rm d}u+\frac{{\rm d}v}{2}-\frac{v-1}{y}{\rm d}y\right)\eta_{1}+\left(\frac{v(v-2)}{y^{2}}{\rm d}u+\frac{{\rm d}v}{2}-\frac{v-1}{y}{\rm d}y\right)\eta_{2}$
$\quad+\left(\frac{2(v-1)}{y^{2}}{\rm d}u-\frac{1}{y}{\rm d}y\right)\eta_{3}\Bigg{]}\,.$ (42)
@@
where $c(u)$ is an arbitrary function of the null coordinate $u$. Notice that we have not yet considered at this point the NMG equations. In the vacuum case, the equations allow for two types of solutions in accordance with the degrees of freedom of the theory Ayón-Beato et al. (2009b). However, in the present case the matter equations restrict the system as above. In particular, notice that the system does not allow for logarithmic modes and so the Breitenlohner-Freedman bound cannot be saturated Breitenlohner and Freedman (1982). Interestingly enough, the YM-CS equations do allow for Schrödinger symmetry to arise Duval et al. (2009). In the equation above this occurs for the special value $k\rho=3$. Lastly, inserting these solutions into the remaining field equations, we find that they are solved by
$\rho^{2}$ $=\frac{\kappa e^{2}-2\mu^{2}}{2\left[\kappa e^{2}\left(k^{2}-\mu^{2}\right)-k^{2}\mu^{2}\right]}\,,$ (44a)
$\Lambda$ $=-\frac{4\mu^{6}\left(5\kappa e^{2}+4k^{2}\right)+4\mu^{4}\left(k^{4}-5\kappa e^{2}k^{2}-\kappa^{2}e^{4}\right)-7\mu^{2}\kappa e^{2}k^{4}+4\kappa^{2}e^{4}k^{4}}{4\mu^{2}\left(\kappa e^{2}-2\mu^{2}\right)^{2}}\,.$ (44b)